Is 2 a solution of 3x+1=6
Answer by oberobic(2304)

No.
.
3x + 1 = 6
.
Subtract 1 from both sides.
.
3x = 5
.
Divide both sides by 3.
.
x = 5/3
.
which can be shown as
x = 1 2/3
.
Always check your work!
.
3x + 1 = ?
3*(5/3) = 5
5 + 1 = 6
Correct!
.
Answer:
x = 5/3, or 1 2/3.
.
Done.